Circular No. 5807

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHAJN-SCHALDACH (1993k) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the 0.9-m Spacewatch telescope at Kitt Peak:

1993 UT  α_{2000}  δ_{2000}  m_{1}
May 27.43324  23^{h}33^{m}53.^{s}03  −1^{o}10′35′′.0
27.44212  23 33 53.67  −1 10 31.4  19.8
27.45291  23 33 54.57  −1 10 26.3  20.0
28.43489  23 35 14.00  −1 03 06.9
28.44249  23 35 14.74  −1 03 03.7  19.6
28.45151  23 35 15.41  −1 02 58.9  19.5

On May 27 there was a coma 10′′ across and a tail extending 0′.55 in p.a. 257^{o}. The indicated correction to the prediction on MPC 16381 (ephemeris on MPC 21959) is ΔT = –0.03 day.

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Further to IAUC 5801, A. Carusi, Consiglio Nazionale delle Ricerche,

Rome (not Rome University), confirms that the orbit on IAUC 5800 shows that the center of the comet’s nuclear train will collide with Jupiter in 1994, and he suggests that — during July 23-27 — this would be the fate of the whole train; the “window” for a collision is in fact more than 30 times the length of the observed train. In the case of a glancing strike, it would be the trailing nuclei (those with a positive change in true anomaly at the 1992 encounter) that survive to escape from jovicentric to heliocentric short-period orbits.

D. K. Yeomans, Jet Propulsion Laboratory, remarks that computations by P. Chodas and himself using observations extending to May 20 indicate that the probability that the center of the nuclear train will collide with Jupiter in late-July 1994 is as high as 64 percent.

NOVA AQUILAE 1993 Further CCD photometry: May 20.047 UT, V = 7.75 ± 0.03, B–V =

+0.55 ± 0.05 (H. Mikuž, Ljubljana, Slovenia; cf. IAUC 5794); 23.028, V = 7.48 ± 0.02 (Mikuž); 25.98, V = 7.96 ± 0.05, B–V = +0.52 ± 0.05, U–B = –0.37 ± 0.1 (D. Hanzl, N. Copernicus Observatory, Brno; comparison star HD 178574, V = 7.56, B–V = +0.48, U–B = –0.02).

1993 May 28 Daniel W. E. Green